5g of a mixture Na2CO3 and Na2SO4 was dissolved in 1000ml of H20. 10 ml of this solution required 8ml of 0.1N H2SO4 solution for neutralisation, the % of Na2CO3 in mixture is?

Grade:12

1 Answers

Vikas TU
7 years ago
N1 V1 = N2 V2
N1 * 10 = 8 * 0.1
N1 = 0.08
M1 = N1/v.f = 0.08/2 = 0.04M of 5g mixture
Total moles = 0.04*1000 = 40mm
moles o f Na2CO3 = x/106and
moles of Na2SO4 = (5-x)/142
40 x 10^-3 = x/106 +  (5-x)/142 solve for x.